### v1

# Differentiable Physics-based System Identification for Robotic Manipulation of Elastoplastic Materials

Abstract
Robotic manipulation of volumetric elastoplastic deformable materials, from foods such as dough to construction materials like clay, is in its infancy, largely due to the difficulty of modelling and perception in a high-dimensional space. Simulating the dynamics of such materials is computationally expensive. It tends to suffer from inaccurately estimated physics parameters of the materials and the environment, impeding high-precision manipulation. Estimating such parameters from raw point clouds captured by optical cameras suffers further from heavy occlusions. To address this challenge, this work introduces a novel Differentiable Physics-based System Identification (DPSI) framework that enables a robot arm to infer the physics parameters of elastoplastic materials and the environment using simple manipulation motions and incomplete 3D point clouds, aligning the simulation with the real world. Extensive experiments show that with only a single real-world interaction, the estimated parameters, Young’s modulus, Poisson’s ratio, yield stress and friction coefficients, can accurately simulate visually and physically realistic deformation behaviours induced by unseen and long-horizon manipulation motions. Additionally, the DPSI framework inherently provides physically intuitive interpretations for the parameters in contrast to black-box approaches such as deep neural networks.
keywords:
Elastoplastic material manipulation, system identification, differentiable physics, deformable material manipulation
1 Introduction
Figure 1: The proposed system identification framework enables a robot to interact with elastoplastic material via simple manipulation motions (orange box) and then identify the physics parameters of the real-world manipulation dynamics. The parameters are found using gradients computed, through differentiable simulation, from a differentiable point-cloud-based similarity function between the real and simulated observations of the manipulated material (blue and red boxes). These parameters then enable accurate simulations that allow the real-world grounding of motion planning, trajectory optimisation or policy learning techniques (cyan box).
Despite the recognised importance of robotic manipulation of deformable materials, this topic remains underexplored, particularly when it comes to high-precision manipulation of volumetric elastoplastic materials. A primary challenge in this area arises from the materials’ infinite degrees of freedom (DoFs), leading to highly unpredictable deformation dynamics.
The intrinsic complexity of these dynamics inhibits the direct application of conventional robotic motion planning methods, which typically require explicit physics models for all concerned objects Latombe (2012). Learning approaches, such as reinforcement learning (RL), often involve training an agent to learn to interpret its perception and take actions through inefficient trial & error in a realistic physics-based simulation Collins et al. (2021); Kroemer et al. (2021), which is both challenging and largely unavailable when it comes to deformable materials.
In contrast to the well-studied rigid body dynamics in robotics, where motions can be predicted and controlled using well-defined equations of motion and deterministic models Featherstone (2014), deformable materials do not follow such straightforward patterns. Directly applying these methods is problematic because it is extremely difficult and often infeasible to accurately model and perceive real-world elastoplastic materials and measure the underlying physics parameters that govern their motions and deformations  Arriola-Rios et al. (2020); Yin et al. (2021). As a result, achieving high-precision manipulation for such materials with motion planning or data-driven techniques is challenging due to the high computational cost and the lack of techniques to capture the dynamics accurately.
To close this gap, this research proposes a Differentiable Physics-based System Identification (DPSI) framework for the robotic manipulation of volumetric elastoplastic materials. Our framework can efficiently estimate key physics parameters governing material deformation dynamics using minimal and simple manipulation motions. The estimated physics parameters enable accurate material simulation for long-horizon predictions of real-world elastoplastic deformation behaviours.
The workflow of the proposed DPSI framework can be summarised as follows. As shown in Figure 1, a robot equipped with an in-hand 3D camera (Zivid) and three end-effectors is deployed to manipulate the elastoplastic object (e.g., play dough). Before manipulation, the robot takes multi-view point clouds of the object to minimise occlusions. These point clouds are used to create the initial particle system for the material point methods (MPM)-based simulation. The robot then performs a manipulation motion on the object and captures point clouds of the deformed state. Physics simulations of the same manipulation are run with the same initial state and motion, whose resultant particle states are compared to the real-world deformed state through variants of the Chamfer distance (CD) and the earth mover’s distance (EMD) loss. The parameters of the simulated physics models are updated after every simulation to minimise the loss. To facilitate fast optimisation, we use a differentiable simulator built on the TaiChi auto-differentiation mechanism, which allows automatic gradient computation from the losses and gradient-based optimisation for the physics parameters Hu et al. (2019, 2020).
Our approach achieves unprecedented simulation-to-real alignment accuracy, characterised by the integration of the following novel features.
High-fidelity physics: Unlike previous methods that either employ non-physics-based models (e.g., neural networks) or highly simplified material geometry representations (e.g., sparse keypoints), we use high-fidelity physics-based simulation powered by the MPM Jiang et al. (2016), which simulates materials as Lagrangian particles and keeps track of their positions and velocities. It achieves faster simulation by computing the motions, deformation gradients, and frictional contacts on a background Eulerian grid Stomakhin et al. (2013); Jiang et al. (2016); Gao et al. (2017); Hu et al. (2018). MPM-based simulations provide highly efficient and realistic simulation with high physical plausibility by closely following real-world physics laws such as Newton’s laws and elastic and plastic energy conservation models.
Incomplete & noisy observations: Unlike existing works that rely on synthetic videos with complete sequences of perfect observations Murthy et al. (2020); Li et al. (2023); Kaneko (2024), our framework uses 3D point clouds to observe real-world object geometries. Capturing the full depth of an object during manipulation is impractical due to occlusions caused by the end-effector or environment. This means that only the point clouds before and after a manipulation motion are practical to obtain and informative enough to observe the full geometry of the deformed object in real-world experiments. In addition, real-world point clouds tend to suffer from inaccurately estimated camera matrices and sensory noises.
Small data, short & simple motions: Using extensive and diverse manipulation motions to collect real-world deformation data is time-consuming and costly. Existing studies demand a significant number of real-world interactions or complete sequences of simulation videos to identify object deformations under various motions, yet still resulting in simulations with insufficient accuracy for real-world applications Lin et al. (2022); Shi et al. (2023); Li et al. (2023); Kaneko (2024). Our goal is to recover physics parameters that enable accurate predictions of long-horizon, unseen and complicated elastoplastic material manipulation dynamics, using minimal simple and short real-world interactions.
Joint parameter estimation: We aim to jointly estimate the physics parameters provided by physics models. Besides Newton’s laws, we employ the fixed corotated elastic energy model Stomakhin et al. (2012), the von Mises plasticity model Jones (2009) and the dynamic friction model in our simulation. These lead to six key parameters: Young’s modulus $E$, Poisson’s ratio $\nu$, yield stress $\boldsymbol{\sigma}_{y}$, material density $\rho$, and the friction coefficients of the table $\eta_{t}$ and end-effectors $\eta_{m}$ (assuming the three end-effectors share the same coefficient). The first four parameters primarily govern the deformation responses, while the last two handle frictional contacts. These parameters are heavily intertwined in governing the behaviours of the manipulated object and there is no intuitive solution to identify one of them without estimating the effects of other parameters. Therefore, we seek to identify these parameters simultaneously.
Differentiable physics: Identifying the physics parameters in their discretised spaces via search or evolutionary algorithms is computationally slow due to the exponentially growing number of possible combinations as the discretisation becomes finer. While gradient-based optimisation methods offer faster convergence toward the minimum, it is infeasible with most physics simulations because many computation steps are not differentiable and these simulators do not support derivative computations. In this work, we explore the feasibility of optimising system parameters using gradients computed by differentiating loss functions through a physics simulator written by DiffTaiChi, a programming language tailored for GPU-accelerated parallel computation and automatic differentiation Hu et al. (2019, 2020). DiffTaichi generates derivative functions for simulation steps via source code transformation that retains arithmetic intensity and parallelism. It uses a memory-efficient tape to record the order of computation kernels for forward simulation and traverses their derivative functions in the backward order to generate gradients through the computation graph. We build DPSI upon DiffTaiChi and explore the feasibility of directly optimising several physics parameters jointly with gradients computed by differentiating point-cloud-based loss functions through the high-fidelity physics simulator.
Substantial experiments demonstrate that our main contribution, DPSI, can achieve highly accurate simulation-to-reality alignment for elastoplastic materials manipulated by unseen, long horizon and complex motions using minimal simple and short interactions, and noisy and incomplete observations. Results show that when multiple solutions and parameter uncertainty exist, DPSI can provide physically intuitive parameter interpretations that can guide further system identification, model improvement, and motion adaptation. Statistics on the computation costs of DPSI indicates promising practical deployment of the DPSI framework.
The rest of the article reviews related literature, presents formally our method and experiment results, and discusses limitations and future directions.
2 Related works
2.1 Deformable object manipulation
Both model-free and model-based approaches have been taken for manipulating deformable materials. Existing model-free methods often lack manipulation precision due to the absence of physics laws that describe the motion and deformation under complex contacts McConachie et al. (2020); Cherubini et al. (2020); Shi et al. (2023, 2024); Shen et al. (2024). Physics-model-based methods, while more accurate, struggle with aligning the simulated dynamics with the real world and often rely on simplified geometric representations for higher computational efficiency, sacrificing manipulation precision Navarro-Alarcon et al. (2016); Yang et al. (2023); Shetab-Bushehri et al. (2023). With or without a model, many of them use simulations that offer a cost-effective way to test manipulation methods and allow the collection of massive data for learning-based approaches Arriola-Rios et al. (2020); Collins et al. (2021); Yin et al. (2021). However, inaccurate dynamics predictions and control precision are unacceptable for many tasks such as surgery, assembly and disassembly. Therefore, akin to the way humans efficiently learn about object and environment physics properties, this work proposes to actively identify system parameters that enable high-precision simulation of real-world manipulation dynamics for volumetric elastoplastic objects.
2.2 Deformable object modelling
Simulators are essential for advancing robotic manipulation, providing a fast, low-cost alternative to real-world testing Featherstone (2014); Collins et al. (2021). In recent years, arguably the most efficient and accurate simulation for 3D deformable objects is achieved by the material point method (MPM) Jiang et al. (2016). Like most well-known methods for simulating continuum matter, such as position-based dynamics (PBD) and smoothed particle hydrodynamics (SPH) Yin et al. (2021), MPM represents the object as Lagrangian particles and keeps track of their positions and velocities. Unlike pure Lagrangian methods, MPM achieves faster simulation by computing the motions, deformation gradients, and frictional contacts on a background Eulerian grid (akin to the finite element method), governed by elastic and plastic energy functions and Newton’s laws in the form of partial differential equations Stomakhin et al. (2013); Jiang et al. (2016); Gao et al. (2017); Hu et al. (2018). MPM has been proven superior to other methods in terms of efficiency and visual effects for objects that undergo large deformations, fractures, and self-collision Hu et al. (2018). Despite its advantages, accurately identifying physics parameters for real-world robotic manipulation of elastoplastic objects remains an open challenge.
2.3 Deformable object system identification
Unlike rigid bodies Jaques et al. (2022); Heiden et al. (2022); Chen et al. (2022), identifying the physics properties of deformable objects is more complex than for rigid bodies, primarily due to the difficulty in measuring key parameters like material properties and friction coefficients Arriola-Rios et al. (2020); Yin et al. (2021). Previous research has mostly focused on linear and planar deformable objects, such as ropes and cloths Sundaresan et al. (2022); Yang et al. (2022); Caporali et al. (2024).
Early works that sought to identify 3D volumetric material properties have much simpler assumptions and scenarios. For instance, one of the earlier works uses gradient-based optimisation to retrieve the stiffness of a spring system representing elastic deformable objects Lloyd et al. (2007), while another work uses an exhaustive search method to find the value of Poisson’s ratio for an elastic form object Güler et al. (2017). They focus on single parameter identification for volumetric elastic deformable objects with reduced Dofs that under-represent the geometries and deformation behaviours of real-life objects.
Recent efforts like GradSim Murthy et al. (2020) have used differentiable physics and rendering to identify five parameters for elastic objects with much higher DoFs from single-view simulation videos, demonstrating the feasibility of differentiable system identification with synthetic videos. Two following works, PAC-NeRF Li et al. (2023) and LPO Kaneko (2024) propose to jointly reconstruct object geometries (shapes, positions and colours) and physical properties by using a voxel neural radiance field Sun et al. (2022) that performs differentiable rendering and allows gradients to be back-propagated from the image space to the Eulerian grid.
Compared to these works that focused on simulations, our study tackles the more challenging task of system identification for real-world objects through robot interactions. Similarly, we also employ the material point method Hu et al. (2018) and DiffTaiChi Hu et al. (2020) for differentiable physics simulation and study the system identification task without known object geometries. However, we aim to align simulations with real-world dynamics with minimal robot interactions, using only incomplete, occluded, and noisy point cloud data.
3 Method
3.1 Overview
The proposed differentiable physics-based system identification (DPSI) framework, as shown in Figure 2, can be divided into the following modules: differentiable dynamics modelling, real-to-sim object and trajectory reconstruction, and the loss functions. This section starts with a formal problem description and elaborates on each of these modules.
Figure 2: The overall workflow of the proposed differential physics-based system identification (DPSI) framework. Modules in each colour are elaborated in individual subsections of the Method section. Green: differentiable dynamics modelling and the physics parameters. Orange: real-to-sim object reconstruction. Blue: real-to-sim trajectory reconstruction. Pink: optimisation and evaluation loss functions.
3.2 Problem statement
We define the physics parameter identification problem for real-world deformable materials as follows. Denote $\mathcal{X}^{real}_{0}$ and $\mathcal{X}^{real}_{T}$ as the observations of a deformable object before and after a manipulation trajectory $\tau^{real}$ of time duration $T$ is applied, denote $\mathcal{S}^{real}_{0}$ as the observation of the real-world end-effector before the manipulation, the stochastic real-world forward dynamics can be written as follows:
$\mathcal{X}^{real}_{T}\sim p^{real}(\mathcal{X}^{real}_{T}|\mathcal{X}^{real}_{0},\boldsymbol{\theta}^{real},\tau^{real},\mathcal{S}^{real}_{0})$
where $\boldsymbol{\theta}^{real}$ is the set of real-world parameters that govern the behaviours of the dynamical process. As it is unlikely to obtain $p^{real}$ and $\boldsymbol{\theta}^{real}$ explicitly, we use an approximate, deterministic and differentiable dynamics model, $f^{sim}$, to simulate such a dynamical process, written as follows:
$\mathcal{X}^{sim}_{T}=f^{sim}(\mathcal{X}^{rec}_{0},\boldsymbol{\theta}^{sim},\tau^{rec},\mathcal{S}^{rec}_{0})$
where $\boldsymbol{\theta}^{sim}$ is the set of adjustable parameters that govern the behaviours of the simulated deformation and contact process, $\mathcal{X}^{rec}_{0}$ is the reconstructed initial observation of the real object, $\mathcal{X}^{sim}_{T}$ is the resultant observation of the reconstructed object after manipulated by the reconstructed motion $\tau^{rec}$, and $\mathcal{S}^{rec}_{0}$ is the initial observation of the simulated end-effector. In our work, we assume the end-effectors are rigid bodies, meaning the observations of the real or simulated end-effectors are coordinates of their frames relative to the world frame, i.e., $\mathcal{S}\in\mathbb{R}^{3}$, which can be readily retrieved from the robot platform or simulator. As we are interested in point cloud observations from the real world and particles in the simulation, all object observations will be 3D point sets, i.e., $\mathcal{X}\in\mathbb{R}^{N\times 3}$ where $N$ is the number of points or particles.
With these notations, the optimisation problem could be formulated as the minimisation of some distance function between the real-world and simulated observations after manipulation. As the parameters are optimised with a dataset of interaction experiences with the real objects, $\mathcal{D}$, the minimisation problem can be written as follows:
$\min_{\boldsymbol{\theta}^{sim}}\ \frac{1}{|\mathcal{D}|}\sum_{\mathcal{D}}d(\mathcal{X}^{real}_{T},\mathcal{X}^{sim}_{T})$ (1)
The rest of this section will discuss the differentiable simulation dynamics $f^{sim}$, the reconstructed object state, $\mathcal{X}^{rec}$, the reconstructed end-effector trajectory, $\tau^{rec}$, and the loss functions, $d(\mathcal{X}^{real}_{T},\mathcal{X}^{sim}_{T})$.
3.3 Differentiable dynamics modelling
Material point method (MPM): This work employs the mean-least-square material point method (MLS-MPM) to simulate the deformable object manipulation dynamics. MPM is a meshless, hybrid computation scheme that enables efficient computation and preserves high physical fidelity for various materials, especially for elastoplastic materials that undergo large deformation, while the MLS-MPM arises from a novel weak form discretisation of the conservation equations and replaces the shape functions in the force computation with MLS approximators, leading to faster and more realistic simulation of sharp separation of particles and two-way coupling with rigid objects that traditional MPM cannot simulate.
We modified the standard procedures in one simulation step of the MLS-MPM as shown below to incorporate end-effector control following Xian et al. (2023). The readers are referred to the original paper for more details about MPM Stomakhin et al. (2013) and MLS-MPM Hu et al. (2018).
  1. 1.
    Compute particle deformation gradient using the MLS approximation equation.
  2. 2.
    Applied plasticity to recompute deformation gradients and particle stress using the elastic energy model
  3. 3.
    Particle to grid. Use the affine particle-in-cell transform Jiang et al. (2015) to transfer the velocities and masses of the Lagrangian particles to the background Eulerian grid nodes. In our implementation, we assume equal volume and mass for the particles.
  4. 4.
    Update end-effector positions given control inputs.
  5. 5.
    Compute grid node momenta and velocities with gravity applied.
  6. 6.
    Signed-distance field (SDF)-based collision detection with rigid objects (end-effector and boundaries) and frictional contact computation.
  7. 7.
    Grid to particle. Use the affine particle-in-cell transform to transfer the velocities and affine coefficients from the grid nodes to the particles. Perform SDF-based collision detection and frictional contact computation again to minimise particle penetration.
  8. 8.
    Update particle positions with the new velocities.
Elastoplasticity: We assume strain elastoplasticity for the studied objects, meaning that the strain-stress relationship of the material is described using the deformation gradient that can be decomposed into elastic and plastic parts , $\mathbf{F}=\mathbf{F}^{E}\mathbf{F}^{P}$, and the first Piola-Kirchoff stress, $\mathbf{P}$, as $\mathbf{P}=\partial\Psi/\partial\mathbf{F}^{E}$, where $\Psi(\mathbf{F}^{E})$ is the elastic energy density function. In this study, we use the fixed corotated model Stomakhin et al. (2012) as the energy function for its robustness under large deformation, although other constitutive models may be used for different needs. The fixed corotated elastic energy density function is
$\Psi(\mathbf{F}^{E})=\mu||\mathbf{F}^{E}-\mathbf{R}^{E}||^{2}_{F}+\frac{\lambda}{2}(J^{E}-1)^{2}$ (2)
where $J^{E}=\text{det}(\mathbf{F}^{E})$, and $\lambda$ and $\mu$ are the Lamé parameters, such that
$\lambda=\frac{E}{2(1+\nu)}$ (3)
$\mu=\frac{E\nu}{(1+\nu)(1-2\nu)}$ (4)
where $E$ is Young’s modulus and $\nu$ is Poisson’s ratio. Thus, the stress can be calculated as
$\mathbf{P}=\frac{\partial\Psi}{\partial\mathbf{F}^{E}}=2\mu(\mathbf{F}^{E}-\mathbf{R}^{E})(\mathbf{F}^{E})^{T}+\lambda(J^{E}-1)J^{E}$ (5)
where $\mathbf{R}^{E}$ is the rotation matrix that comes from the polar decomposition of the deformation gradient, $\mathbf{F}^{E}=\mathbf{R}^{E}\mathbf{S}^{E}$, where $\mathbf{S}^{E}$ is a symmetric matrix. In practice, they are computed from the singular value decomposition of $\mathbf{F}^{E}$ as $\mathbf{F}^{E}=\mathbf{U}^{E}\mathbf{\Sigma}^{E}(\mathbf{V}^{E})^{T}=\mathbf{U}^{E}(\mathbf{V}^{E})^{T}\mathbf{V}^{E}\mathbf{\Sigma}^{E}(\mathbf{V}^{E})^{T}$, and then $\mathbf{R}^{E}=\mathbf{U}^{E}(\mathbf{V}^{E})^{T}$ and $\mathbf{S}^{E}=\mathbf{V}^{E}\mathbf{\Sigma}^{E}(\mathbf{V}^{E})^{T}$ Baker (2005).
For the $n$-th simulation step, we compute a trial elastic deformation gradient, $\tilde{\mathbf{F}^{E,n}}$. If no plastic deformation appears, meaning $\mathbf{F}^{P,n}=\mathbf{F}^{P,n-1}$, then we assign $\mathbf{F}^{E,n}=\tilde{\mathbf{F}}^{E,n}$. However, if the yield criterion is violated, $\tilde{\mathbf{F}}^{E,n}$ needs to be modified according to the plasticity model. This process of modifying $\tilde{\mathbf{F}}^{E,n}$ is called the return mapping process, which determines how the object would respond plastically.
In this work, we follow Gao et al. (2017) to use the von Mises model Jones (2009) to compute the return mapping, which takes on the associative plastic flow assumption. The projection process of the trial stress outside of the yield region can be described succinctly as $\tilde{\boldsymbol{\epsilon}}^{E,n}-\boldsymbol{\epsilon}^{E,n}=\delta_{\gamma}\tilde{\boldsymbol{s}}^{E,n}/||\tilde{\boldsymbol{s}}^{E,n}||$, where $\tilde{\boldsymbol{\epsilon}}^{E,n}=\log(\tilde{\mathbf{F}}^{E,n})$ and $\boldsymbol{\epsilon}^{E,n}=\log(\mathbf{F}^{E,n})$ are the trial and modified Hencky strains, $\tilde{\boldsymbol{s}}^{E,n}=\tilde{\boldsymbol{\epsilon}}^{E,n}-\ tr(\tilde{\boldsymbol{\epsilon}}^{E,n})/3\ \boldsymbol{I}$ is the trial deviatoric stress, and $\delta_{\gamma}\tilde{\boldsymbol{s}}^{E,n}/||\tilde{\boldsymbol{s}}^{E,n}||$ is the modifying term based on the solution given in Gao et al. (2017). In particular, based on the von Mises yield criterion, $\sqrt{3J_{2}}-\boldsymbol{\sigma}_{y}\leq 0$, where $\boldsymbol{\sigma}_{y}$ is the yield stress and $J_{2}$ is the second deviatoric stress invariant, it was derived that $\delta_{\gamma}=||\tilde{\boldsymbol{s}}^{E,n}||-\boldsymbol{\sigma}_{y}/2\mu$. With the fact that the singular vectors of the trial elastic deformation gradient do not change, the return mapping operation is done on the eigenvalues of the trial Hencky stress, thus on the eigenvalues of the trial elastic deformation gradient. Then, we can obtain the result of the return mapping by taking the exponential of the modified deviatoric stress: $\Sigma^{E,n}=e^{\boldsymbol{\epsilon}^{E,n}}$ and then the modified deformation gradient: $\mathbf{F}^{E,n}=\tilde{\mathbf{U}}^{E,n}\mathbf{\Sigma}^{E}(\tilde{\mathbf{V}}^{E,n})^{T}$.
Note that the Hencky strain formation is only used in plastic response computation. At each simulation step, we compute the trial deformation gradient using the MLS approximation equation, apply plastic response, and then compute the new stress with Equation 5.
Frictional contacts: Collision detection is done by checking the distance of the particles to the surface of the rigid objects using pre-computed signed distance fields (SDFs). For each rigid object (table, and three end-effectors in our case), we employ the procedure described in Park et al. (2019) to generate SDFs for watertight meshes. We assume that the frictional contacts only happen in two cases: between the particles and the table, or between the particles and the end-effector. We also assume uniform friction coefficient distribution over the contact surface of the table and the end-effectors. When a particle or grid node and an object are in contact, we follow Stomakhin et al. (2013) to determine the velocity of the particle and grid nodes using dynamic friction with sticky impulse.
Specifically, for each particle, we calculate the local normal $\overrightarrow{n}$ and the velocity $v_{obj}$ of the rigid object at the particle position. Then, we project the particle velocity $v$ into the object reference frame to have $v_{rel}=v-v_{obj}$. No collision is applied if the particle is separating from the object, i.e., $v_{n}=v_{rel}\cdot\overrightarrow{n}\geq 0$. Define the tangential portion of the relative particle velocity as $v_{t}=v_{rel}-\overrightarrow{n}v_{n}$. For sticky impulse, we set the relative particle velocity $v_{rel}^{\prime}$ after the collision to 00 if $||v_{t}||\leq-\mu v_{n}$, where $\mu$ is the friction coefficient. If the sticky impulse is overcome, dynamics friction is applied, such that $v_{rel}^{\prime}=v_{t}(1+\mu v_{n}/||v_{t}||)$. After the collision, the new particle velocity in the world frame is calculated as $v^{\prime}=v_{rel}^{\prime}+v_{obj}$. We represent the friction coefficients for the table and the manipulators separately as $\mu_{t}$ and $\mu_{m}$.
Differentiable programming: Several programming tools are available for creating differentiable simulations, such as PyTorch Sundaresan et al. (2022); Arnavaz et al. (2023), DiffTaiChi Hu et al. (2020), Jax Schoenholz and Cubuk (2020) and neural-network-based simulators Heiden et al. (2021). We build our simulator based on DiffTaiChi due to its automated differentiation mechanism, GPU-accelerated parallel computation, fast computation kernel evaluation, intuitive Python APIs, rich community support and various promising applications Huang et al. (2021); Lin et al. (2022); Xian et al. (2023).
In particular, for each numerical computation step, DiffTaiChi flattens its computation branches (e.g., boundary and collisions) and replaces mutable local variables with extra local storage variables, producing straight-line codes without mutable variables. It then uses standard source code transform to generate the derivative function based on the adjoint method. To compute gradients with a loss function, DiffTaiChi records the order of computation kernels and the scalar variables in the forward simulation direction, and then it computes the gradients for the concerned variables by evaluating the derivative functions in the reversed simulation direction Hu et al. (2020). We build a simulator based on this programming language to allow automatic gradient computation for the physics parameters. The codes for the simulator will be open-sourced upon paper acceptance.
In summary, the elastic and plastic models describe the deformation behaviours of the object using Young’s modulus $E$, Poisson’s ratio $\nu$ and yield stress $\sigma_{y}$, the computation of the particle and grid velocities will involve another parameter, the object density $\rho$, and the frictional contact processes are controlled by respective friction coefficients: table friction coefficient $\mu_{t}$ and manipulator friction coefficient $\mu_{m}$. We optimise these six parameters to align the simulation to the real-world dynamics.
3.4 Real-to-sim object reconstruction
Real-world platform: To collect real-world data, we set up a deformable object manipulation system where a Kuka IIWA LBR 14 industrial arm Kuka (2024) is equipped with a Zivid One+ medium camera Zivid (2024) for perceiving the object, and one of the three end-effectors, namely a rectangular cuboid, a cylinder roller and a bullet-shaped object for collecting interaction experiences with different contact geometries, as shown at the bottom-left of Figure 1. We use plasticine (non-hardening modelling clay) as the main material for our experiments.
Real-world perception: We create a multi-view point-cloud capture and fusion process to obtain the real-world observation of the deformable object $\mathcal{X}^{real}_{0}$ and $\mathcal{X}^{real}_{T}$. In particular, for each object state, the robot arm moves to six poses around the object for the camera to take a point cloud of the scene. The point clouds are fused and then cropped to contain only the points of the object.
These observations are noisy and incomplete in three senses. Firstly, due to camera calibration error, the fusion result always exhibits a ±$\sim\pm 3$ mm discrepancy. Secondly, due to joint limits, the robot arm cannot reach poses that allow the camera to capture the bottom part of the object. To simplify the problem, we assume that the angle between the object boundary and the table surface is equal to or greater than $90^{\circ}$, which allows the camera to capture as much as possible the bottom part of the object. We then project all points to its bottom to form a closed surface. Note that the initial configuration of the object can be manually shaped to satisfy this assumption of contact angle but the end configuration of the object after being manipulated is out of manual control, which tends to have more occlusions. Lastly, we only take the observations before and after applying the manipulation motions, without providing the intermediate observations during the manipulation, because it is impractical to do so when too much occlusion occurs during manipulation.
Reconstruction pipeline: To simulate the object as a set of particles, we design a pipeline to reconstruct the object particle system $\mathcal{X}^{rec}$ from the given fusion point cloud, as shown by the orange boxes in Figure 2. Firstly, we down-sample the processed fusion point cloud and create a watertight mesh using the ball-pivoting (BP) algorithm Bernardini et al. (1999). The artificial bottom surface points prevent the BP algorithm from creating a non-flat bottom mesh surface. To sample particles, we create a set of voxels that fill in the reconstructed mesh surface uniformly given a resolution ($1080\times 1080$). Lastly, we fill particles into the voxels with a given density ($4\times 10^{6}$ particles$/m^{3}$). We assume that the particles are uniformly distributed within the mesh.
For the end-effectors, we assume they are rigid bodies. In simulation, we keep track of the coordinate of its frame and a pre-computed SDF for collision detection.
Optimisation motions (short and simple) Duration (sec.) Real Num. waypoints Sim. Num. waypoints
Lv. 1 Poking-1 $-z$ $0.015$ m, $+z$ $0.03$ m 0.86 38 87
Poking-2 $-z$ $0.02$ m, $+z$ $0.03$ m 0.93 40 94
Lv. 2 Poking-shifting-1 $-z$ $0.02$ m, $-x$ $0.03$ m, $+z$ $0.03$ m 1.52 68 152
Poking-shifting-2 $-z$ $0.02$ m, $+x$ $0.03$ m, $+z$ $0.03$ m 1.50 54 153
Validation motions (unseen, longer, more complex contacts)
Flattening (cylinder) $-z$ $0.025$ m, $+x$ $0.025$ m, $+z$ $0.025$ m, $-x$ $0.025$ m, $-z$ $0.025$ m, $-x$ $0.025$ m, $+z$ $0.025$ m 3.74 275 378
Triple-poking (round) $+y$ $0.025$ m, $-z$ $0.025$ m, $+z$ $0.025$ m, $-y$ $0.025$ m, $-z$ $0.025$ m, $+z$ $0.025$ m, $-y$ $0.025$ m, $-z$ $0.025$ m, $+z$ $0.025$ m 4.55 328 460
Poking-180-rotating (rectangle) $-z$ $0.025$ m, about $+z$ $180^{\circ}$, $+z$ $0.025$ m 6.23 504 625
Table 1: Waypoint designs and statistics of the interaction motions. For collecting optimisation data, there are two motions for each of the two levels of contact complexity. For out-of-distribution validation, one motion per end-effector is designed. Directions of the waypoints are relative to the robot base frame, as shown in Figure 1. Each waypoint of the real trajectories takes uneven time interval, while each waypoint in simulation takes exactly $0.01$ second.
3.5 Real-to-sim trajectory reconstruction
Manipulation motions: As we are interested in recovering real-world dynamics with small data and short, simple motions, we design experiments in two contact complexity levels, for which we create two poking motions and two poking-shifting motions for optimisation and in-distribution validation, and three longer motions for out-of-distribution validation. These motions are summarised in Table 1 in terms of consecutive trajectory segments in the world frame.
The first contact level focuses on identifying the parameters that primarily govern the deformation of the object (Young’s modulus, Poisson’s ratio, yield stress and material density) using two poking motions, which press the object down by a certain distance. The second level further includes the friction coefficients, using two poking-shifting motions, which press down the object and make horizontal shifting movements. For out-of-distribution validation, three longer motions with more drastic contact processes are created. Each motion uses a different end-effector. The triple-poking motion is designed for the round end-effector to validate the long-term deformation prediction with small fraction influences. The flattening motion is designed to validate the long-term deformation and frictional contact prediction under large linear movements, using the cylinder end-effector. The pressing-180-rotating motion is designed to validate the long-term deformation and frictional contact prediction under large rotational movements using the rectangle end-effector.
All motions start from a configuration where the end-effector tip is positioned at the top centre of the object. ROS Quigley et al. (2009) and the MoveIt!-based Görner et al. (2019) OMPL planner Sucan et al. (2012) are used to plan real-world motion trajectories $\tau^{real}$. For calculating each motion plan, we pass a series of waypoints to the MoveIt! planner by discretising each segment of the motion with an interval of $0.002$ m or $5^{\circ}$.
A challenging phenomenon that occurs during our data collection process is that the object tends to stick to the end-effectors after contact is made. If we allow this to happen and assume that the object always drops down eventually, the optimisation process will be extremely difficult due to a large uncertain dynamics process of the object dropping from the air. For example, the object may bounce out of the workspace or even off the table. This further exacerbates the difficulty of optimising with only the start and end observations of the object. However, it is beyond the capability of current perception hardware and thus a future research direction. To simplify the problem, we take a simple workaround by covering the end-effector with a thin layer of flour before each manipulation motion is executed. This greatly prevents the sticking phenomenon from happening. Note that the end-effector friction coefficient we are optimising for is then the one covered by flour instead of the original value.
Time duration constrained reconstruction: To simulate end-effector motions, the real-world trajectory generated by MoveIt! is inconvenient as it has uneven time differences between consecutive waypoints and the simulation can only handle a constant step size $dt$. Thus, for a motion segment between a pair of waypoints, we reconstruct its simulation counterpart to have constant velocity by dividing the travelled distance by the real time duration which is provided by the motion planner, and then we discretise the segment with a constant $dt$. In this study, we set $dt=0.01$ second for better simulation stability. A larger $dt$ will result in a too-high compounding error during simulation stepping, while a too small $dt$ will demand too much computation. The statistics of the real and reconstructed motions are summarised in Table 1. It can be seen that the validation motions are much longer than the optimisation ones. The MoveIt! trajectories and the reconstructed ones are saved as .npy files and will be open to the public upon acceptance.
3.6 Loss functions
We use four loss functions to calculate the difference between the simulated and real object states. As we are dealing with points and particles, it is natural to select point-based metrics. Therefore, we employ the two most common distance metrics for point sets, namely, the Chamfer distance (CD) and the earth mover’s distance (EMD). Given two point sets $\mathcal{X}_{0}$ and $\mathcal{X}_{1}$, the CD and EMD are defined as follows.
$\begin{split}d_{CD}(\mathcal{X}_{0},\mathcal{X}_{1})=&\frac{1}{|\mathcal{X}_{0}|}\sum_{\boldsymbol{x}\in\mathcal{X}_{0}}\min_{\boldsymbol{y}\in\mathcal{X}_{1}}||\boldsymbol{x}-\boldsymbol{y}||_{2}+\\ &\frac{1}{|\mathcal{X}_{1}|}\sum_{\boldsymbol{y}\in\mathcal{X}_{1}}\min_{\boldsymbol{x}\in\mathcal{X}_{0}}||\boldsymbol{x}-\boldsymbol{y}||_{2}\end{split}$
$d_{EMD}(\mathcal{X}_{0},\mathcal{X}_{1})=\frac{1}{|\mathcal{X}_{0}|}\sum_{\boldsymbol{x}\in\mathcal{X}_{0}}\min_{\phi^{EMD}:\mathcal{X}_{0}\rightarrow\mathcal{X}_{1}}||\boldsymbol{x}-\phi(\boldsymbol{x})||_{2}$
where $\boldsymbol{x}$ and $\boldsymbol{y}$ denote the 3D coordinates of the points, $\phi^{EMD}$ denotes a one-to-one injective mapping that only exists when $|\mathcal{X}_{0}|\leq|\mathcal{X}_{1}|$. In practice, a linear assignment algorithm is used to calculate $\phi^{EMD}$. We also find that the three average operations reduce the magnitudes of the gradients of the physics parameters to an order of $4$ to $5$, which is undesired for optimisation; therefore, we use the distances without averaging the values over the point sets.
To calculate the loss between the real and simulated end configurations, one can use the original fusion point cloud $\mathcal{X}_{T}^{real}$ or the reconstructed particle systems $\mathcal{X}_{T}^{rec}$ as the target observations. Thus, there are two variants for each loss: $d_{PCD-CD}(\mathcal{X}_{T}^{real},\mathcal{X}_{T}^{sim})$, $d_{PRT-CD}(\mathcal{X}_{T}^{rec},\mathcal{X}_{T}^{sim})$, $d_{PCD-EMD}(\mathcal{X}_{T}^{real},\mathcal{X}_{T}^{sim})$, $d_{PRT-EMD}(\mathcal{X}_{T}^{rec},\mathcal{X}_{T}^{sim})$. When reconstructing $\mathcal{X}_{T}^{rec}$, we decrease the particle density for filling the voxels slowly until $|\mathcal{X}_{T}^{rec}|\leq|\mathcal{X}_{T}^{sim}|$ is satisfied. When computing losses involving the real-world point clouds, we downsample it with a voxel radius of $0.005$ so that the number of points in it is about $\frac{1}{6}$ of the number of particles in the reconstructed system. This is because a point cloud only contains points on the surface, while the reconstructed particle system is non-hollow.
Finally, although the CD and EMD losses are very common in calculating point set distance, they are not intuitive to visualise and understand. Also, as revealed by the experiment results shown in Figures 4 and  5, the CD and EMD losses focus on different spatial aspects of two point sets, which make them biased for result analysis. To make results easier to analyse and compare in an unbiased way, we further calculate the heightmaps of the fusion point cloud and the simulated particles, denoted as $\mathcal{I}_{T}^{real}$ and $\mathcal{I}_{T}^{sim}$. Specifically, we discretise a $0.11\times 0.11\text{m}^{2}$ square area centred at the centre of the object into a $32\times 32$ grid. For each cell of the grid, we find the point or particle in the x-y range with the highest $z$ value and assign that value as the height of that cell. To compare the distance between two heightmaps, we simply take their summed pixel difference as follows.
$d_{Heightmap}=\sum_{i}\sum_{j}||\mathcal{I}_{T,ij}^{real}-\mathcal{I}_{T,ij}^{sim}||_{2}$
In practice, we use the fusion point cloud without downsampling to compute $\mathcal{I}_{T}^{real}$. However, a particle density that is too low will result in void cells when generating $\mathcal{I}_{T}^{sim}$, and we found that a density value of $4\times 10^{6}$ particles per cubic metre is good enough for generating $32\times 32$ heightmaps. The heightmap distance is only used for evaluation purposes because we found it leads to very poor optimisation results when used as a loss function.
4 Results
This section presents the design of the experiment and the result analyses including the optimisation and generalisation performances of the DPSI framework in identifying the physical characteristics of elastoplastic matter through simple robot manipulation.
4.1 Experiment design
The performance of the DPSI framework is examined in three steps: loss landscape analysis, in-distribution performance analysis and out-of-distribution generalisation analysis. The first step visualises the four loss functions (PCD-CD, PRT-CD, PCD-EMD, and PRT-EMD) to help understand how the physics parameters are related to the loss value distributions. The second step investigates the in-distribution performance of the proposed DPSI framework, i.e., whether it can produce realistic simulations for unseen object configurations with the same motions used in parameter identification. The loss landscape and in-distribution performance analyses are carried out at two levels of contact complexity. At each level, the performances are compared across four loss functions and five optimisation datasets (see data collection description below). The last step explores the out-of-distribution generalisability of the DPSI by inspecting the simulation accuracy of three unseen longer motions using identified parameters from the contact complexity level 2 experiment.
Data collection: For each optimisation motion (i.e., Poking-1, Poking-2, Poking-shifting-1, Poking-shifting-2), we collect datapoints by executing the motions on the objects with different end-effectors. Each datapoint contains the multi-viewpoint point clouds of the object before and after a manipulation motion. For each of the two complexity levels, five optimisation datasets are created to examine the data hungriness of the proposed DPSI framework.
With level (Lv.) 1 as an example, the first dataset is created with both motions (Poking-1 and Poking-2). Each motion is performed twice using three different end-effectors. Therefore, $6$ datapoints are created for each motion, and a total of $12$ datapoints are created for both motions. The second dataset for Lv. 1 is similar, except that only $1$ datapoint is created from each motion-effector pair, resulting in $3$ datapoints for each motion, and a total of $6$ datapoints for both motions. The other three datasets for Lv. 1 are more straightforward, with only one datapoint collected for each dataset. The three datapoints are created using the three end-effectors individually, performing the second motion from Lv. 1 (i.e. Poking-2). We name these datasets as 12-mix, 6-mix, 1-rectangle, 1-round, and 1-cylinder, respectively. Data collection for Lv. 2 is similar, except that the two motions of Poking-1 and Poking-2 are replaced by Poking-shifting-1 and Poking-shifting-2, respectively. The datasets will be open-sourced upon the acceptance of the article.
For in-distribution validation, with each contact complexity level, we collect two extra datapoints from the second motion (i.e., Poking-2 and Poking-shifting-2) with all three end-effectors, resulting in $12$ validation datapoints. For out-of-distribution validation, where we collect two datapoints with each of the three long-horizon motions, resulting in $6$ datapoints.
For each datapoint, the real-world object is initially roughly shaped into a convex shape. We then acquire its point cloud by fusing the captured multi-view data. Next, we calculate the coordinates of the top centre of the object, where the end-effector tip will be moved to, and the motion will always be executed from there. We perform the same point cloud capturing and fusion process after a motion is executed. Each datapoint takes about $2$ to $3$ minutes to collect, with most of the time spent on capturing point clouds.
Optimisation & validation: At each contact level, for each pair of loss function and dataset, gradient descent is carried out with the Adam algorithm Kingma and Ba (2015) with three random seeds for $100$ iterations (gradient updates). For each datapoint, the simulation loads the initial particle configuration, simulates the motion and produces the resultant particle state, which is used to compute losses and gradients. Each optimisation iteration goes over all datapoints within a dataset and takes the average gradients to update the parameters. In-distribution validation is done after every gradient update with the validation dataset, simulating all datapoints and calculating the losses.
Table 2 displays the step sizes and the value ranges of the parameters of interest for gradient updates. To determine the appropriate step sizes, we randomly sample $20$ sets of parameter values and compute the statistical means and standard deviations of the gradients for each parameter based on the 12-mix dataset. We then randomly initialise the parameters, run the Adam optimiser for $100$ gradient updates with uniformly randomly generated gradients using the statistical means and deviations with different step sizes, and select the step size that allows the randomised parameters to converge in about $70$ gradient updates. The value ranges of the parameters are selected according to either the definitions of the physics models Jones (2009); Stomakhin et al. (2013) or values reported by various research on manual parameter identification experiments with similar real-world materials, such as soil and clay StructX (2014-2024); Waheed and Asmael (2023). However, a too-high value of Young’s modulus $E$ was found to cause numerical instability in the simulation with the particles exploding away. Thus, an empirical upper bound for $E$ was determined by slowly increasing it until the simulation became unstable.
Parameter $E$ (kPa) $\nu$ $\rho$ (kg/m3) $\boldsymbol{\sigma}_{y}$ (kPa) $\eta_{t}$ $\eta_{m}$
Step size $4000$ $0.01$ $10$ $500$ $0.01$ $0.01$
Min value $10000$ $0.01$ $1000$ $1000$ $0.01$ $0.01$
Max value $300000$ $0.48$ $2000$ $20000$ $2.0$ $2.0$
Table 2: step sizes and the ranges of values for the parameters: Young’s modulus $E$, Poisson’s ratio $\nu$, yield stress $\boldsymbol{\sigma}_{y}$, material density $\rho$, and the friction coefficients of the table $\eta_{t}$ and end-effector $\eta_{m}$.
4.2 Loss landscape analysis
The loss landscapes, including point-cloud and particle chamfer distance (PCD CD, PRT CD), point-cloud and particle earth mover’s distance (PCD EMD, PRT EMD), are computed over three pairs of physics parameters: $(E,\nu)$, $(\boldsymbol{\sigma}_{y},\rho)$ and $(\mu_{g},\mu_{e})$. For each pair, we set other parameters to fixed values and compute the losses with the parameters of interest discretised into $30$ intervals. Figure 3 displays the loss landscapes computed at the two contact complexity levels and five datasets. To facilitate distribution-level comparison, the loss landscapes are centralised to have zero means (subtracting the mean of the matrix). The following observations can be drawn.
Figure 3: Loss landscapes (centralised to have zero means) at level-1 (left) and level-2 (right) contact complexity over pairs of physics parameters. Darker colours represent higher loss values. Parameters: Young’s modulus ($E$) against Poisson ratio ($\nu$), yield stress ($\boldsymbol{\sigma}_{y}$) against material density ($\rho$) for both levels, and table friction coefficient ($\mu_{t}$) against manipulator friction coefficient ($\mu_{m}$) for level 2. Other parameter values are fixed while computing the losses over one pair of parameters. Each column presents the losses computed with a different dataset. Each row presents the values of a different loss function. PCD: point cloud. PRT: particle system. CD: Chamfer distance. EMD: earth mover’s distance.
Firstly, we compare the landscapes in Figure 3 vertically to examine the sensitivity of the loss functions against different parameters. At contact complexity level 1, it shows that the CD and EMD losses exhibit quite similar value changing directions along the $E$, $\nu$ and $\rho$ axes, while quite the opposite directions along the $\boldsymbol{\sigma}_{y}$ axis. This can be observed from the vertically reversed colour distributions between the first and last two rows of landscapes against $\boldsymbol{\sigma}_{y}$ and $\rho$. On the other hand, the losses at level 2 contact complexity exhibit similar distributional patterns along all parameter axes. In addition, at level-2 contact complexity, the PRT CD loss is more sensitive (exhibiting more drastic value changes) along the $\nu$ axis than other losses (see the second row of the level-2 landscapes against $E$ and $\nu$). In contrast, the values of the PRT CD and both EMD losses are more sensitive along the $\boldsymbol{\sigma}_{y}$ axis (the last three rows of the level-2 landscapes against $\boldsymbol{\sigma}_{y}$ and $\rho$). The level-2 landscapes against the friction coefficients ($\mu_{t}$ and $\mu_{m}$) show that all losses distribute very similarly over these two parameters. Overall, the EMD losses seem to be more sensitive to all the parameters, as their loss landscapes tend to be less flat.
These observations indicate that, with the adopted von Mises plasticity model, whose plastic deformations are governed by the yield stress ($\boldsymbol{\sigma}_{y}$) of the material, the CD and EMD losses may focus on different aspects of the point sets, as they tend to disagree with each other distributionally along the $\boldsymbol{\sigma}_{y}$ axis. Also, we may expect the EMD losses to perform better in parameter identification because they exhibit higher sensitivity against most parameters.
Secondly, by comparing the landscapes horizontally, Figure 3 shows the distributional patterns of most (not all) loss landscapes are quite similar across the five datasets. This means that the loss values, and thus the optimisation processes, are not sensitive to the number of datapoints used in computation. This observation allows us to expect the recovery of physics parameters with small data, even with a single interaction experience.
Thirdly, many of the loss distributions display large areas of flat regions, where the loss values are very similar. This indicates many optimisation saddle points and that multiple parameter value combinations may serve as plausible solutions, describing the same physical characteristics. Whether this is an issue from an inaccurate physics model or a general fact of the real-world dynamics remains to be determined.
4.3 In-distribution performances
This subsection investigates the in-distribution performance of the parameter identification task through gradient descent at two levels of contact complexity. For each level, experiments are conducted with the four loss functions and five datasets. To evaluate the performances thoroughly, the four loss functions and a heightmap-based distance function are used to evaluate the differences between the real and simulated manipulation results using the in-distribution validation dataset.This enables the observation of the influences on other distance functions from minimising each objective. In particular, this subsection investigates the following questions.
  • Can the loss functions be minimised (does optimisation converge)?
  • Do the loss functions agree with each other in such system identification tasks?
  • If local minima appear and multiple solutions exist, do they produce visually distinct manipulation results?
  • How does the number of datapoints affect optimisation?
  • Does DPSI produce parameter values that are physically realistic and interpretable?
Quantitative results: We start by analysing the quantitative results of the parameter identification task. The changes of the validation losses and parameter values over the course of optimisation at both contact complexity levels are shown in Figure 4 and 5. The best parameter values corresponding to the lowest validation heightmap loss among the three random seeds are summarised in Table 3.
First of all, we start by observing the tendency of convergence. The top five rows in Figure 4 and 5 show that, at both contact complexity levels, most of the individual validation loss curves (dotted lines) tend to stabilise and converge, which indicates that DPSI can effectively converge to local minima. The last four rows in Figure 4 and the last six rows in Figure 5 reveal that the parameters converge to different solutions. Table 3 also shows that multiple parameter solutions exist for a low validation heightmap distance at both levels of experiments. These observations mean the parameter identification task at both contact complexity levels does converge but has multiple local minima and solutions, aligned with the large flat regions observed from the loss landscapes.
Figure 4: Validation losses (top five rows) and parameter values (last four rows) over $100$ gradient updates at level-1 contact complexity. Each column presents the results of optimising with a different dataset. Each row shows the changes of an evaluation metric or a parameter, denoted on the left. In each figure, different colours indicate the results of minimising a different loss function, as labelled by the legend on the top. For the top five rows, each colour has three dotted lines corresponding to the results of three random seeds and a solid line corresponding to their means. For the last four rows, each colour has three solid lines corresponding to the results of three random seeds.
However, the parameter distributions in Figure 5 clearly show that some parameters converge to much smaller and distinct value regions than those found in the level-1 experiments. More specifically, the values of Young’s modulus ($E$), Poisson’s ratio ($\nu$) and yield stress ($\boldsymbol{\sigma}_{y}$) become more converged, while they are more dispersed in the level-1 results. Interestingly, the material density ($\rho$) remains quite dispersed at both contact complexity levels. This is not surprising because, compared to the level-1 motions, the higher contact complexity induced by the shifting parts of the level-2 motions naturally poses a stricter selection range for the physics parameter values. Also, some parameter values are closer to empirical studies of soft/hard/saturated clay in the level-2 experiments. For example, the Young’s modulus values are closer to the reported range of $5000$ to $54000$ kPa, the material density seem to be closer to $1400$ kg/m3, and Poisson’s ratio are closer to the reported range of $0.4$ to $0.5$ Waheed and Asmael (2023); StructX (2014-2024). However, the reported yield stress values for clays ($210$ to $600$ kPa) Rehman et al. (2018) were found to make the material collapse in our simulator. A comparison of simulating the materials using the reported parameters in the literature and the ones found by DPSI is procided in the fifth subsections of the Results section.
Figure 5: Validation losses (top five rows) and parameter values (last six rows) over $100$ gradient updates at level-2 contact complexity. Each column presents the results of optimising with a different dataset. Each row shows the changes of an evaluation metric or a parameter, denoted on the left. In each figure, different colours indicate the results of minimising a different loss function, as labelled by the legend on the top. For the top five rows, each colour has three dotted lines corresponding to the results of three random seeds and a solid line corresponding to their means. For the last four rows, each colour has three solid lines corresponding to the results of three random seeds.
Secondly, the top five rows in Figure 4 reveal that, in most cases, the CD and EMD losses have a negative correlation. For example, the red curves of the first column show that minimising the PRT EMD loss reduces both EMD losses and the heightmap distance, but increases both CD losses. Also, Table 3 shows that optimising the CD losses tends to produce large yield stress values ($\boldsymbol{\sigma}_{y}$), while optimising the EMD losses tends to produce small ones. Similar observations can be made from the level-2 results, but only the PCD CD validation losses tend to exhibit opposite tendencies compared to the other losses (observe the blue lines that deviate from others in Figure 5). This is aligned with the analyses of the loss landscapes at both contact complexity levels, where reversed distributional patterns of the CD and EMD losses were observed. Furthermore, the EMD losses tend to produce smaller validation heightmap distances at both levels as shown by the green and red lines in Figure 4 and 5 and the average heightmap distances in Table 3, indicating better sim2real alignment. This again is aligned with the analysis of the level-2 loss landscapes, where only the PCD CD loss distributes significantly differently along the $\boldsymbol{\sigma}_{y}$ axis.
Dataset
Contact complexity Loss Metric/Parameter 12-mix 6-mix 1-rec. 1-round 1-cyl. Avg. Height map (mm)
Min. height map (mm) 2267.44 2176.21 1973.81 2183.30 2327.01 2185.55
PCD Young’s modulus (kPa) 300000 10000 300000 300000 294991
CD Poisson’s ratio 0.480 0.480 0.480 0.480 0.348
Yield stress (kPa) 20000 12367 1000 7527 20000
Material density (kg/m3) 1000 1380 1225 1000 1000
Min. height map (mm) 2287.21 2293.42 2285.98 2311.87 2290.71 2273.83
PRT Young’s modulus (kPa) 189900 300000 179373 196393 20335
CD Poisson’s ratio 0.165 0.010 0.223 0.010 0.178
Yield stress (kPa) 7373 10286 7016 12358 2466
Level Material density (kg/m3) 1157 1088 1210 1000 1797
1 Min. height map (mm) 2024.78 2039.95 2040.51 2186.90 2272.50 2112.93
PCD Young’s modulus (kPa) 265420 237337 270940 300000 193790
EMD Poisson’s ratio 0.445 0.439 0.419 0.480 0.436
Yield stress (kPa) 1403 1036 1382 5839 10138
Material density (kg/m3) 1487 1379 1482 1000 1009
Min. height map (mm) 2000.36 1999.14 1980.35 1999.34 2016.10 1999.06
PRT Young’s modulus (kPa) 300000 300000 300000 300000 271247
EMD Poisson’s ratio 0.453 0.480 0.480 0.478 0.442
Yield stress (kPa) 1000 1000 1059 1000 1000
Material density (kg/m3) 1542 1670 1214 1518 2000
Min. height map (mm) 3163.64 3091.73 3622.58 3423.00 3125.04 3285.20
PCD Young’s modulus (kPa) 47700 42136 91785 132840 28679
CD Poisson’s ratio 0.444 0.451 0.364 0.095 0.403
Yield stress (kPa) 11407 10479 7839 15590 14514
Material density (kg/m3) 1810 1919 1819 1000 1880
Table friction 0.725 0.696 1.816 0.696 0.715
Manipulator friction 1.495 1.448 1.264 1.037 0.704
Min. height map (mm) 3088.31 2865.14 3104.15 3031.56 2999.36 3017.70
PRT Young’s modulus (kPa) 66514 47786 28023 24068 256786
CD Poisson’s ratio 0.331 0.376 0.317 0.309 0.451
Yield stress (kPa) 1904 1809 8363 8063 1537
Material density (kg/m3) 1267 1284 1926 1965 1225
Table friction 0.473 0.488 0.414 0.419 0.487
Level Manipulator friction 0.494 0.604 0.249 0.293 0.084
2 Min. height map (mm) 2933.23 2667.43 2959.43 2895.74 2979.04 2886.97
PCD Young’s modulus (kPa) 48459 14959 135207 113474 12017
EMD Poisson’s ratio 0.392 0.400 0.374 0.339 0.369
Yield stress (kPa) 1417 1176 1359 1340 10921
Material density (kg/m3) 1319 1409 1313 1331 1995
Table friction 0.544 0.525 0.508 0.453 0.434
Manipulator friction 0.805 1.308 0.194 0.381 0.878
Min. height map (mm) 2598.28 3060.10 2847.69 2760.83 3172.49 2887.88
PRT Young’s modulus (kPa) 27779 10000 222856 144827 10000
EMD Poisson’s ratio 0.359 0.330 0.435 0.383 0.303
Yield stress (kPa) 1008 9635 1000 1000 7377
Material density (kg/m3) 1396 2000 1492 1445 2000
Table friction 0.487 0.373 0.311 0.543 0.411
Manipulator friction 1.013 1.478 0.015 0.625 1.433
Table 3: The parameter values corresponding to the lowest validation heightmap distance in each optimisation case at both contact complexity levels. Bolded texts denote the three lowest losses at each level.
Thirdly, the results show that it is possible to obtain comparable performances with just one datapoint at both contact complexity levels. This can be concluded by comparing the results horizontally: 1) most validation curves of the top five rows in Figs. 4 and 5 with 1-datapoint show highly similar tendencies with the 6-mix and 12-mix datasets, and 2) the last four rows in Figs. 4 and last six rows in Figure 5 also show that the parameters found with different datasets mostly converge to similar value regions. Moreover, Table 3 shows that the lowest validation heightmap distance at level-1 contact complexity and the third lowest at level-2 are achieved by optimising with the 1-rec. and 1-round datasets. These results are aligned with the analysis of the loss landscapes, where similar loss distributions are observed among different datasets.
Figure 6: Validation losses (top five rows) and parameter values (last four rows) over $100$ gradient updates at level-1 contact complexity from optimising the PRT EMD loss function with an extra set of $8$ random seeds. Each column presents the results of optimising with a different dataset. Each row shows the changes of an evaluation metric or a parameter, denoted on the left. In each figure, different colours indicate the results of a different random seed, as labelled by the legend on the top.
Lastly, we perform extra experiments of optimising the PCD EMD loss with the three 1-datapoint datasets to examine the random seed sensitivity of DPSI. The validation losses and parameters are shown in Figure 6. On the one hand, these runs all converge to stable (though different) loss and parameter values, still reveal the nagative relationship between the CD and EMD losses, and can achieve comparable performances with results from larger datasets. On the other hand, it can be seen that the main cause of converging to different solutions is the different initial parameter values, which were determined by the peusdo random process controlled by different random seed values. This is not suprising as gradient-descent algorithms with complex physics problems are known to be sensitive to the choice of the initial solution Hu et al. (2020); Antonova et al. (2023). Therefore, with a good initial guess of the solution, DPSI is likely to achieve desirable system identification results.
Figure 7: Visualisations of two configurations of the simulated objects and their heightmaps after applying the second motions at contact level 1 (left) and 2 (right), simulated with the best set of physics parameters. Darker colours of the heightmaps indicate greater heights of the object. For each combination of the loss function and dataset at each level, three trajectories are simulated (3 effectors $\times$ 1 datapoints). The results for each trajectory are grouped with the ground truth placed on the left. In each group, a row shows the results corresponding to a loss function and a column results corresponding to a dataset.
Qualitative results: We now examine the manipulation processes simulated using the best physics parameters corresponding to the lowest validation heightmap losses as shown in Table 3. Figure 7 shows the particles and heightmaps after the manipulation for two object configurations for each pair of contact complexity and end-effector.
Firstly, all the simulated particles and heightmaps post-manipulation are highly similar to the ground truths across the loss functions and datasets. This shows that minimising any of the four loss functions with any of the five datasets can indeed reproduce visually plausible manipulation results close to the real-world system. It indicates that DPSI is not data-hungry and it is robust to the choice of common point-based loss objectives. These characteristics are highly preferred in robotic applications, as they lead to smaller data collection costs, simpler observation preprocessing, and simpler loss function engineering.
Secondly, by examining the details of the heightmaps more carefully, it shows that optimising the CD losses tends to produce heightmaps with darker colours (which represent greater heights of the object), while optimising the EMD losses tends to produce heightmaps where the objects look slightly bigger and wider. It indicates that optimising the CD losses tends to produce particles that match the height of the ground truth shapes, i.e., in the $z$ direction, while optimising the EMD losses tends to match the $x$ and $y$ directions. This is especially true at level-1 contact complexity. At level 2, the PRT CD loss sometimes (compare the second rows of the level-2 heightmaps with the rectangle and round end-effectors) changes its focus onto the $x$ and $y$ directions, producing lighter-colour heightmaps.
This may be caused by the fact that these losses distribute differently along the yield stress axis (the CD and EMD losses at level 1, the PRT CD loss at level 2). According to physics intuitions, a smaller yield stress will cause the object to yield more easily and respond more drastically to the poking forces, hence the more spreading in $x$ and $y$ directions and more compressing in the $z$ direction, and vice versa. One can see that a greater yield stress value in Table 3 corresponds to a darker heightmap in Figure 7, and vice versa.
Thirdly, though the particles and heightmaps are very similar, they are produced by quite different parameter combinations. For example, Table 3 reveals a positive correlation between the yield stress ($\boldsymbol{\sigma}_{y}$) and material density ($\rho$). One can see that, although the simulated results are similar, a larger $\boldsymbol{\sigma}_{y}$ is always accompanied by a larger $\rho$, and vice versa. This is realistic because when the object becomes heavier (greater $\rho$), it needs to be more difficult to yield; otherwise, it will collapse due to gravity. If the object is light (lower $\rho$), a smaller yield stress value allows it to exhibit the returning effect after compression.
In addition, from the level-2 results in Figure 7, one can observe another correlation between the friction coefficients and the material density. They are the three key parameters that determine how much the object will be moved in the shifting motion direction. The parameters in Table 3 and their corresponding visualisations show that,
  1. 1.
    with similar weights ($\rho$), the object is moved at a longer distance when the manipulator frictions ($\mu_{m}$) is greater (compare the parameters and visualisations between the results from the PCD EMD loss with the 12-mix and 1-rec. datasets)
  2. 2.
    with similar manipulator frictions ($\mu_{m}$), the object is moved further when the table friction ($\mu_{t}$) is low (compare the parameters and visualisations between the PCD CD loss with the 1-round dataset and the PRT EMD loss with the 12-mix dataset)
  3. 3.
    with similar friction values, the object is moved further when it is lighter (compare the results from the PCD EMD loss with the 12-mix and the 1-cyl. datasets)
These relationships between parameters and visualisations align with our understanding of real-world physics, demonstrating that DPSI can produce physically plausible and interpretable parameter values.
Fourthly, the visualisations at both levels also reveal limitations and potential improvement directions of the physics model. By comparing with the ground truths, the simulated contact areas of the objects always deform more sharply with insufficient elastic returning. The real objects, on the other hand, tend to respond more elastically after being plastically deformed, hence the higher and smoother surfaces shown in the ground truths. This is more obvious for contact areas with sharp edges such as the four edges of the rectangle end-effector or the sharp sides of the cylinder end-effector. In addition, at level-2 cases, with the effectors shifting horizontally, some simulated particles are unrealistically displaced and stay floating. In the real world, the displaced parts would fall because of gravity. What’s more, the shifting motion in the real world causes the whole object to tilt in the moving direction, while in the simulation the contact impact tends to remain in a much smaller region around the contact area. These inconsistencies can be caused not only by inaccurate modelling, but also by various computation approximations, such as time integration method, simulation step size, contact handling, etc.
Summary: In short, this subsection presents a detailed examination of the in-distribution performances of DPSI with motions at two levels of contact complexity. By analysing the validation loss curves, the parameter values and their manipulation visualisations, we can answer the proposed questions:
  • The loss functions can be minimised, demonstrating the feasibility of DPSI even in the presence of noisy and incomplete point cloud observations.
  • The loss functions produce similar visualisations, but do not always agree with each other on the found parameter values. The CD and EMD losses seem to focus on quite different spatial aspects of the point sets. They distribute differently along the yield stress ($\boldsymbol{\sigma}_{y}$) axis and produce different yield stress values. However, it seems that, as the optimisation motion becomes more complex, the loss functions agree more on the parameter solutions.
  • There are many local minima and possible parameter solutions, but they produce visually and physically similar manipulation results.
  • The number of datapoints has a minor influence on the optimisation performances both quantitatively and qualitatively, indicating that DPSI is not data-hungry even in the presence of real-world perception challenges.
  • Discussions on the visualisations and parameter values show that DPSI can produce physics parameter values with physically realistic and interpretations.
4.4 Generalisation
This subsection will look at the out-of-distribution performance of DPSI by visualising the manipulation processes of three unseen and much longer motions that induce more complex contact dynamics with the best parameters (Table 3) found at the level-2 contact complexity identification task.
Figure 8: Visualising the real long-horizon manipulation motions and their simulations based on the recovered parameters. Every four rows display the real and the top-three best simulations of a motion-object pair.
To recall, there are three unseen motions. The triple-poking motion moves the round end-effector to poke the object at three different locations along the $y$ axis. The flattening motion moves the cylinder end-effector to press and flatten the object towards the positive and negative $x$ directions. The poking-180-rotating motion moves the rectangle end-effector to press the object down and then rotate $180$ degrees about the $z$ axis. The triple-poking motion is less relevant to the friction coefficients, while the last two motions involve complex frictional contacts. Table 5 summarises the heightmap distances of the simulated results of the unseen motions (each with two object configurations). Figure 8 visualises the trajectories of the real and the top three best simulation trajectories. The animation videos of them are included in Supplementary materials. Figure 9 presents the resultant particles and their heightmaps of each motion and object configuration. The following observations can be drawn.
First, in Figure 8, (as well as the animation video), it can be seen that, for all motion and object configuration pairs, the real and simulated trajectories are highly similar. Figure 9 also shows that, although discrepancies exist, the simulated particles and their heightmaps of the manipulation results are also highly similar across the loss functions and datasets. This visual similarity proves that DPSI can indeed generalise to unseen, longer, and more complex manipulation motions, exhibiting robustness against perception and data challenges. Moreover, we note that it is the single-datapoint cases that achieve two-thirds of the top-three lowest heightmap losses across the three unseen motions as shown in Table 5.
Secondly, for the unseen and more complex motions, the found physics parameters still retain their physical meanings and interpretability. Evidence supporting this can be found by observing the differences between the heightmaps in Figure 9 produced by different parameters.
  • The first example is that, when the yield stress ($\boldsymbol{\sigma}_{y}$) is smaller, the resultant heightmaps of the unseen motions tend to be lighter, meaning the objects are more compressed down.
  • Another example is that, when moving the cylinder end-effector along the $x$ axis after pressing down the object, a higher manipulator friction coefficient ($\eta_{m}$) causes larger whole-body displacement of the object, while a smaller value causes the effector to slip away and compress down the particles along its moving direction. This is shown more clearly in the bottom-right case (cylinder end-effector) of Figure 9, where the resultant object bodies that are more concentrated to the heightmap centres correspond to larger manipulator friction coefficients, while the objects whose upper part of the heightmaps are flattened correspond to smaller one.
  • The last example is that, for the poking-180-rotating motion results shown at the top of Figure 9, when the objects are more rotated by the end-effector, at least two of the following three conditions can be verified true from Table 3 at the same time: a relatively small material density, a relatively small table friction coefficient, and a relatively large manipulator friction coefficient. This is physically intuitive as any two of the three conditions would make the object easier to move.
Thirdly, the found solutions are not globally optimal. Despite the visual similarity, Table 5 shows that the best generalisation results are not always achieved by the cases with the best in-distribution validation performances, as the underlined and bolded texts rarely coincide. Qualitatively, from the heightmaps and trajectories, one can also recognise certain visual discrepancies between the real objects and simulations. This indicates a certain degree of discrepancy between the found parameters and the actual solutions. Therefore, it shows that the data (and the motions that produce them) used for optimisation in our experiments are insufficient. Although multiple causes could be possible, it is out of the scope of this work to design better data collection processes for finding the optimal parameters.
Finally, the modelling limitations observed in the in-distribution visualisation still exist in the unseen motion visualisations (Figs. 8 and 9). The insufficient elastic returning at contact areas especially with sharp manipulator edges can be seen from all visualisations where the primary contact areas of the object rarely return to heights close to the real object. Floating displaced particles can be observed from the poking-180-rotating motion visualisations where many particles are carried away during rotations. The insufficient spreading of the influence of contacts can be recognised from the first object configuration of the flattening motion visualisations where the object is not tilted enough in the second half of the trajectory.
In summary, the results and discussions above strongly support the argument that DPSI is indeed capable of predicting unseen, longer, and more complex deformable object manipulation dynamics by optimising with very few real-world noisy and incomplete point-cloud data collected by simple and short interacting motions.
Motion Particle density $4\text{x}10^{7}$ $6\text{x}10^{7}$ $8\text{x}10^{7}$
Number of particles 1946 2808 3926
CP. Forward time (s) 9.56 9.83 9.23
Forward time (s) 1.45 1.61 2.17
Poking-1 Forward GPU (GB) $\sim$1.6 $\sim$1.8 $\sim$1.9
CP. Backward time (s) 24.69 24.71 24.42
Backward time (s) 2.27 2.39 2.56
Backward GPU (GB) $\sim$1.7 $\sim$1.8 $\sim$1.9
$100$ FB-Iterations (min) 6.71 7.18 8.37
Number of particles 3880 5836 7738
CP. Forward time (s) 11.09 14.63 23.97
Poking- Forward time (s) 2.87 7.06 16.78
shifting-1 Forward GPU (GB) $\sim$1.9 $\sim$2.2 $\sim$2.5
CP. Backward time (s) 26.15 27.67 27.34
Backward time (s) 4.31 5.42 5.79
Backward GPU (GB) $\sim$1.9 $\sim$2.2 $\sim$2.5
$100$ FB-iterations (min) 12.47 21.30 38.10
Number of particles 3059 4688 6271
CP. Forward time (s) 14.02 15.19 15.94
Forward time (s) 6.08 6.54 7.88
Flattening Forward GPU (GB) $\sim$1.8 $\sim$1.9 $\sim$2.1
CP. Backward time (s) 31.95 34.83 34.93
Backward time (s) 10.14 12.29 13.03
Backward GPU (GB) $\sim$1.8 $\sim$1.9 $\sim$2.1
$100$ FB-iterations (min) 27.53 31.90 35.35
Table 4: Computational costs (time and GPU memory comsuption) of the TaiChi-based differentiable simulator used in the DPSI framework with a Nvidia GeForce RX 4090 GPU. The costs of simulations with three motions and particle densities are reported. CP.: compile-time costs. The first time a program is run takes extra time to compile TaiChi kernels. $100$ FB-iterations (min): the runtime of $100$ forward and backward computations in minutes.
Motion Object Num. datapoint 12 6 1 (rec.) 1 (round) 1 (cyl.)
PCD CD 6178.32 6089.17 6418.81 6620.95 6146.67
Config. PRT CD 6238.82 6256.54 6166.92 6167.83 5964.65
0 PCD EMD 6105.29 5997.06 6034.08 6042.02 6157.98
Triple- PRT EMD 6499.63 6201.23 6193.93 6478.28 6243.00
poking PCD CD 9860.21 10189.94 10288.10 10215.02 10085.14
Config. PRT CD 10247.12 10267.20 10323.14 10228.00 $9$.92
1 PCD EMD 10378.19 10539.97 10300.59 10345.66 9768.11
PRT EMD 10715.33 9735.01 10374.79 10320.68 9903.31
PCD CD 5241.38 5270.79 4777.82 5015.08 5186.37
Config. PRT CD 5241.58 5145.21 5306.65 5388.45 5100.66
0 PCD EMD 5465.86 4698.51 5678.32 5658.92 5405.66
Flattening PRT EMD 5667.67 5108.31 6190.99 5887.79 5208.29
PCD CD 10890.47 10452.40 9600.17 9688.61 11074.55
Config. PRT CD 11897.35 11450.01 10934.94 10692.00 12251.92
1 PCD EMD 10985.88 10353.19 12189.07 12421.18 10476.08
PRT EMD 10836.55 10977.26 12123.76 11286.20 10990.40
PCD CD 5161.34 5085.54 5302.23 7091.04 4979.64
Config. PRT CD 6900.43 6342.54 5024.69 5018.13 8079.13
0 PCD EMD 6321.36 5288.02 6467.52 7036.70 4981.49
Poking-180 PRT EMD 5447.25 5051.74 7483.40 5844.29 5188.73
rotating PCD CD 9436.61 10015.51 10034.32 9212.14 10396.27
Config. PRT CD 8577.18 8613.41 10461.23 10385.83 8268.34
1 PCD EMD 8585.16 9126.36 8745.00 8401.15 9929.82
PRT EMD 8800.82 9394.39 9102.43 8765.50 9583.31
Table 5: The lowest pixel-wise heightmap distances (in mm) achieved by the three unseen motions (each with two object configurations), simulated by the best parameters found at the level 2 contact complexity identification task. Bolded texts denote the top three lowest losses. Underlined texts denote the losses achieved by the cases which have achieved the top-three lowest in-distribution validation performances.
Figure 9: Visualisations of the particles and their heightmaps after applying the unseen motions, simulated with the best set of physics parameters found at the level 2 contact complexity experiments. Darker colours of the heightmaps indicate greater heights of the object. For each motion, there are two object configurations. The results for each trajectory are grouped with the ground truth placed on the left. In each group, a row shows the results corresponding to a loss function and a column results corresponding to a dataset.
4.5 Remark on computation cost
In this subsection, we report the running time and GPU memory needed for forward and backward computations for the DPSI framework with different particle densities and motions. The statistics are summarised in Table 4, which were obtained using a Nvidia GeForce RTX 4090 GPU.
It can be seen that the computation time and memory increase as the number of simulation particles and the length of motion increase. Because of the pre-compilation of kernels when they are first called in a TaiChi programme, the first forward and backward passes always take much longer and the later repetitive calls are much faster, as denoted by ”CP. Foward time (s)”.
With a paritcle numbers around $2000$ to $4000$, the runtimes of an DPSI optimisation run that consists of $100$ iterations of forward-backward computations using $1$ datapoint can be done in about $10$ minutes. As the shown by Figures 4 and 5, most runs converge within $50$ iterations, which means that a plausible solutions can be found by DPSI in about $5$ minutes. With a good initial guess of the parameters, as shown by Figure LABEL: (to be updated), this runtime can be further reduced. In addition, the memory comsuptions for these simulations are within $3$ GB, which further permits implementations on portible and small GPUs. Therefore, this indicates that DPSI meets the computation costs criteria for practical deployment.
5 Conclusion
This work addresses the important problem of closing the gap between simulated and real-world manipulations of elastoplastic objects. In particular, we propose a differentiable physics-based system identification framework (DPSI) that can identify physics parameters through gradient descent algorithms. Substantial experiments demonstrate that the proposed framework can identify parameters that reproduce quantitatively and qualitatively realistic elastoplastic object manipulation dynamics in the presence of real-world perceptual and data collection challenges. These challenges include 1) simple and short motions, 2) incomplete trajectories, 3) object occlusions, 4) point cloud noises, and 5) small data. The proposed framework is the first example of system identification with differentiable physics-based, particle-based simulation for robotic volumetric elastoplastic object manipulation. It serves as the foundation for faster and more accurate real-world deployments of deformable object manipulation.
In addition, with the use of physics-based dynamics models, the identified parameters are physically meaningful. The experiments reveal that different parameter values found through optimisation can be interpreted in a way that is aligned with our understanding and intuitions about real-world physics. Therefore, the DPSI framework not only gives users confidence in the simulation controlled by the physics models and these parameters, but also provides users with intuitive angles to identify the limitations of the reconstructed manipulation dynamics.
5.1 Limitations and future research
Several limitations can be observed from the experiments conducted in this work. The most obvious would be the under-representativeness of the physics models that describe the underlying manipulation dynamics. As discussed in the Results section, there is a lack of elastic returning effect at the areas of contact with sharp edges of objects, the insufficient spread of force impact that causes the deformation to stay near the local contact area, and the artefacts of floating displaced particles. Though these inaccuracies may be negligible for coarse manipulations that appear more often in our daily lives, they would lead to unacceptable solutions for high-precision manipulation tasks such as surgery, soft object assembly, etc. In addition, various approximations in the MLS-MPM and programming implementations also contribute to the simulation inaccuracy. As such, future research is needed to develop more accurate physics models, numerical approximations, and coding implementations to simulate object deformations.
Secondly, an important assumption of this work is that the deformation behaviour of the target material in need of parameter identification can be simulated by the selected elastoplasticity model. Thus, it should be noted that when the target material’s behaviour is largely underrepresented by the selected physics model, DPSI would not be able to produce realistic simulations. In other words, the manual selection of the elastoplasticity model is of high importance for the proposed DPSI framework. For example, the use of the von Mises plasticity model will make it difficult for DPSI to recover parameter values that accurately reflect realistic soil or foam behaviours, which are better captured by the Drucker–Prager model.
Thirdly, an important lesson learnt from the results is that increased complexity of motion and contact mode lead to less local minima and better system identification accuracy. The number of local minima indicates that the collected data is insufficient to fully induce the correct values of the concerned parameters. Thus, in the future, we envision a better framework that incorporates learning-based approaches to allow the automatic selection of diverse tools and interaction motions to achieve more efficient and accurate parameter identification.
Another limitation comes from the means of capturing real-world observations. In this work, we employ the multi-view fusion point clouds as the observation space, which suffers from noise and the inaccurate estimate of the camera extrinsic matrix. In the future, new methods should be developed to reduce point cloud noises and improve the precision of camera calibration. Lastly, the simulated scene in this work is limited to an end-effector, a table surface, and a target elastoplastic object. Efforts are needed to extend the framework to support more realistic and complex contact dynamics. More experiments are also needed to examine the feasibility of large-scale simulation in terms of efficiency and accuracy.
ACKNOWLEDGMENT
This work was supported by the Engineering and Physical Sciences Research Council (grant No. EP/X018962/1).